Solve the following equations using logarithms- show all your work. Round to 3 decimal places.
2x = 30
5x - 1 = 3x
3.53x + 1 = 65.4
16x - 4 = 33 - x
7x - 2 = 5x

Answer by Alan3354(69443)

Solve the following equations using logarithms- show all your work. Round to 3 decimal places.
2x = 30
x = 15
-----------------
5x - 1 = 3x
2x = 1
x = 1/2
-------------------
3.53x + 1 = 65.4
16x - 4 = 33 - x
-----------------
7x - 2 = 5x
2x = 2
x = 1
---------------
I don't know why logs would be used.